OpenStudy (jojo12g):

What is the solution to the equation? 3(x – 5) = 8x – 30 A. x = –5 B. x = 5 C. x = –3 D. x = 3

OpenStudy (aveline):

Use the distributive property: 3x-15=8x-30 add 15 to both sides: 3x=8x-15 subtract 8x -5x=-15 divide by -5 x=3

OpenStudy (jojo12g):

thank you so much